OpenStudy (anonymous):

Find an exact solution. 1 + tan(θ)/sin(θ)= 0

OpenStudy (tkhunny):

Are we forgetting out Order of Operations? Do you mean \((1+\tan(\theta))/\sin(\theta) = 0\)?

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

Okay, what's your plan? You need a plan.

OpenStudy (tkhunny):

\(\dfrac{1+\tan(\theta)}{sin(\theta)} = \dfrac{1}{\sin(\theta)} + \dfrac{\tan(\theta)}{\sin(\theta)}\)... I'm just tinkering. Anything pop out at you?

OpenStudy (anonymous):

That would equal sec(x) + csc(x)=0 right?

OpenStudy (tkhunny):

I suppose, but does that get us anywhere?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."